OpenStudy (anonymous):

Classify the solutions of (1)/(x+4)+1/2=1/(x+4) as extraneous or non-extraneous.

OpenStudy (anonymous):

Choices x = -4; extraneous x = -4; non-extraneous x = -8; extraneous x = -8; non-extraneous

OpenStudy (anonymous):

\[\frac{ 1 }{ x+4 }+\frac{ 1 }{ 2 }=\frac{ 1 }{ x+4 }\]

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@AnitaS.